# Two new approaches to multiple canonical correlation analysis for repeated measures data

### v1

@@ 1. Introduction
The multiple CCA problem was first considered by Horst [24] whose solution is called the “maximum correlation method”. It, however, suffers from severe problems with convergence, [8]. Kettenring [29] proposed a different solution (also for $T=1$) to this problem, and named it the SUMCOR. Also relevant to our work, [10, 28, 27] extended the CCA to the case of several data sets. Their method is known as the “generalized canonical correlation analysis (GCCA)” or the “multiple-set canonical correlation analysis (MCCA)”. It has been considered by many authors, e.g., [15, 14, 33, 40, 48, 51, 50, 13, 55, 52].
@@
We adapt the methodology of multiple CCA to time-dependent observations of the multiple features of interest, that were obtained from different experimental units. Our first approach relies on kernel embeddings of each block into Reproducing Kernel Hilbert spaces. For this methodology, which covers multiple CCA for many other data structures, we derive consistency rates that cover data scenarios, in which the cross-covariance operators are not compact or the observations in the sample have some dependence structure among the experimental units. Neither case, to the best of our knowledge, has been studied, but are relevant for data scenarios that motivate our work. In fact, assuming the opposite, i.e. independence and compactness, is hard to justify.
Our second approach is an extension of multiple CCA to functional data indexed by time, which we think is particularly suitable for the problems that motivate this work. In the functional context, we view the repeated measurements as smooth time courses of indicators, and show that the consistency rates derived for the multiple kernel CCA carry over to this extension. To illustrate the functional context, in the competitiveness index example, suppose $Y_{l,j}^{(k)}(t)$ denotes the value of indicator $j$ in the feature $l$ for country $k$ in year $t$. Unless a country experiences a dramatic change, like a government overthrow or a financial crash, it is reasonable to assume that the values of the $Y_{l,j}^{(k)}(t)$ evolve smoothly from year to year. This will be true for other examples, like the performance of sectors of the economy or agriculture. In such cases, a regularization of the time trajectories might lead to better-performing procedures.
@@ 2. Multiple kernel canonical correlation analysis in the case of repeated measures data
In this section, we introduce multiple canonical correlation analysis and the corresponding estimators based on repeated measures data. We show that the consistency results in CCA can be extended to multiple CCA, including cases, where
@@
Instead the function space related representation (9), the above finite-dimensional representation is suitable for numerical solvers. Now, we formulate the assumptions for our multiple kernel canonical correlations consistency result. Our theory is based on the proof strategies of [12], but we relax several of their assumptions.
@@
As explained previously, some regularization is required, because the covariance operator estimates do not have full rank, and thus, are not invertible. For our asymptotic considerations, the regularization sequence must converge to 0, but not too fast.
@@
While Assumptions 1-3 are fairly standard, Assumption 4 is more delicate. Below, we present two examples, in which Assumption 4 is satisfied. Example 1 highlights that it is possible to drop the requirement of compact cross-covariance operators, which is imposed throughout the literature. Example 2 is focused on incorporating dependence between the experimental units. In fact, this is more realistic than assuming independence for most real data settings.
@@
Applying a similar argument with respect to the $\mathrm{L}^{2}$-norm to the other summands entails the same rate for them. Hence, $\mathbb{E}(\|\boldsymbol{\widehat{C}}_{1,2}-\boldsymbol{C}_{1,2}\|_{\mathrm{op}})=O(1/\sqrt{n})$. Thus, it is possible to not assume compactness of the cross-covariance operators $\boldsymbol{C}_{i,j}$ if the sample, respectively, the embeddings of the sample are bounded. Non-compact cross-covariance operators can actually arise from very simple dependence structures, e.g. $Z_{1}^{(i)}=Z_{2}^{(i)}$, $i\in\{1,\ldots,N\}$. The latter is a very trivial example in the context of CCA, but useful to point out that there exist more complex dependence structures that cannot be described by compact cross-covariance operators. This was observed in [12] who provided a sufficient condition for a cross-covariance operator to be Hilbert-Schmidt. However, this condition is not satisfied in general.
@@
holds, because the covariance operators are bounded and invertible, which implies boundedness of their inverse square-root. By setting $\boldsymbol{A}=\boldsymbol{C}_{i,i}+\epsilon_{n}\cdot\boldsymbol{I}$ and $\boldsymbol{B}=\boldsymbol{C}_{i,i}$ in
@@ 3. Multiple functional CCA
and to evluate them for each datum in the sample and attempt to cluster subjects by their values. This is typically done by drawing a scatterplot on the $(U^{(i)},U^{(j)})$ plane. Subjects with similar similar $(U^{(i)},U^{(j)})$ values are considered as sharing important features.
@@
Let the random variables $Y_{l}[k]$ take values in a separable Hilbert space, have finite second moments and Assumptions 2-4 hold. Moreover, the conditions on the eigenvalues and operators imposed in Theroem 1 are satisfied. Then, the results of Theorem 1 directly carry over to multiple functional CCA described in this section.
@@ 4.1.1. Multivariate repeated measures data
Figure 3. Multivariate repeated measures data for 16 voivodeships and seven regions in the system of the first two multiple kernel canonical variables $(U^{(1)},U^{(2)})$
@@ 4.1.2. Multivariate functional data
We used a Fourier basis with nine components ($B_{ij}=9,\ i\in\{1,\ldots,L\},\ j\in\{1,\ldots,p_{i}\}$) to express the agriculture data as functional data. Using the transformed data, multiple functional canonical correlation analysis was performed (Table 1). The multivariate functional data for the 16 voivodeships in the system of the first two multiple functional canonical variables $(U^{(1)},U^{(2)})$ are shown in Figure 5 which shows that in the system of multiple canonical variables, the voivodeships are grouped into compact clusters belonging to specific macroregions. Additionally, looking at canonical correlations (Table 1), we see that higher values were obtained for the multivariate functional approach. Thus, for these data, multiple functional canonical variables have stronger discrimination ability than the multiple kernel canonical variables.
Figure 5. Multivariate functional data for 16 voivodeships and seven regions in the system of the first two multiple functional canonical variables $(U^{(1)},U^{(2)})$
@@ 12. Innovation 5
Figure 7. Multivariate functional data for GCI dataset (115 countries divided into five groups) in the system of the first two multiple functional canonical variables $(U^{(1)},U^{(2)})$ (with 95% confidence normal ellipses)

### v2

# Two approaches to multiple canonical correlation analysis for repeated measures data

@@ 1. Introduction
The multiple CCA problem was first considered by Horst [24] whose solution is called the “maximum correlation method”. It, however, suffers from severe problems with numerical convergence, [8]. Kettenring [29] proposed a different solution (also for $T=1$) to this problem, and named it the SUMCOR. Also relevant to our work, [10, 28, 27] extended the CCA to the case of several data sets. Their method is known as the “generalized canonical correlation analysis (GCCA)” or the “multiple-set canonical correlation analysis (MCCA)”. It has been considered by many authors, e.g., [15, 14, 33, 40, 48, 51, 50, 13, 55, 52].
@@
We extend the methodology of multiple CCA to the new setting of time-dependent observations of the multiple features of interest obtained from different experimental units. We propose two approaches. Our first approach relies on kernel embeddings of each block into a Reproducing Kernel Hilbert space. We derive consistency rates that cover data scenarios in which the cross-covariance operators are not compact or the observations in the sample have some dependence structure among the experimental units. Neither case, to the best of our knowledge, has been studied, but they are relevant for data scenarios that motivate our work. Assuming independence and compactness is often an assumption of convenience, so our theory has practical relevance. Our second approach is an extension of multiple CCA to functional data indexed by time, which we think is particularly suitable for the problems that motivate this work. In the functional context, we view the repeated measurements as smooth time courses. We show that the consistency rates derived for the multiple kernel CCA carry over to this extension. To illustrate the functional context, in the competitiveness index example, suppose $Y_{l,j}^{(k)}(t)$ denotes the value of indicator $j$ in the feature $l$ for country $k$ in year $t$. Unless a country experiences a dramatic change, like a government overthrow or a financial crash, it is reasonable to assume that the values of the $Y_{l,j}^{(k)}(t)$ evolve smoothly from year to year. This will be true for other examples, like the performance of sectors of the economy or agriculture. In such cases, a regularization of the time trajectories might lead to better-performing procedures.
Unlike in regression, no specific response variables exist in the data structures we study. Our methodology does not rely on assigning the roles of predictor and response variables. The consistency of multiple CCA has not yet been theoretically analyzed in the literature.
Our framework provides a unified view of several extensions of CCA, allowing both finite-dimensional and functional data to be treated within the same operator-theoretic setting. This perspective facilitates a rigorous comparison between kernel- and function-based approaches. The proposed approaches may serve as a foundation for advances in consistent inference for high-dimensional dependent data and regularized and sparse variants of multiple CCA.
@@ 2. Multiple kernel canonical correlation analysis in the case of repeated measures data
In this section, we introduce multiple canonical correlation analysis and the corresponding estimators based on repeated measures data. We show that the consistency results in CCA can be extended to multiple CCA, including cases where
@@
Instead, the function space related representation (9), the above finite-dimensional representation is suitable for numerical solvers. Now, we formulate the assumptions for our multiple kernel canonical correlations consistency result. Our theory is based on the proof strategies of [12], but we relax several of their assumptions.
@@
As explained previously, some regularization is required because the covariance operator estimates do not have full rank and thus are not invertible. For our asymptotic considerations, the regularization sequence must converge to 0, but not too fast.
@@
While Assumptions 1-3 are fairly standard, Assumption 4 is more delicate. Below, we present two examples in which Assumption 4 is satisfied. Example 1 highlights that it is possible to drop the requirement of compact cross-covariance operators, which is imposed throughout the literature. Example 2 is focused on incorporating dependence between the experimental units. In fact, this is more realistic than assuming independence for most real data settings.
@@
Applying a similar argument with respect to the $\mathrm{L}^{2}$-norm to the other summands entails the same rate for them. Hence, $\mathbb{E}(\|\boldsymbol{\widehat{C}}_{1,2}-\boldsymbol{C}_{1,2}\|_{\mathrm{op}})=O(1/\sqrt{n})$. Thus, it is possible to not assume compactness of the cross-covariance operators $\boldsymbol{C}_{i,j}$ if the sample, respectively, the embeddings of the sample are bounded. Non-compact cross-covariance operators can actually arise from very simple dependence structures, e.g. $Z_{1}^{(i)}=Z_{2}^{(i)}$, $i\in\{1,\ldots,N\}$. The latter is a very trivial example in the context of CCA, but useful to point out that there exist more complex dependence structures that cannot be described by compact cross-covariance operators. This was observed in [12], who provided a sufficient condition for a cross-covariance operator to be Hilbert-Schmidt. However, this condition is not satisfied in general.
@@
holds, because the covariance operators are bounded and invertible, which implies boundedness of their inverse square root. By setting $\boldsymbol{A}=\boldsymbol{C}_{i,i}+\epsilon_{n}\cdot\boldsymbol{I}$ and $\boldsymbol{B}=\boldsymbol{C}_{i,i}$ in
@@ 3. Multiple functional CCA
and to evaluate them for each datum in the sample and attempt to cluster subjects by their values. This is typically done by drawing a scatterplot on the $(U^{(i)},U^{(j)})$ plane. Subjects with similar similar $(U^{(i)},U^{(j)})$ values are considered as sharing important features.
@@
Let the random variables $Y_{l}[k]$ take values in a separable Hilbert space, have finite second moments, and Assumptions 2-4 hold. Moreover, the conditions on the eigenvalues and operators imposed in Theorem 1 are satisfied. Then, the results of Theorem 1 directly carry over to multiple functional CCA described in this section.
@@ 4.1.1. Multivariate repeated measures data
Figure 3. Scatterplot for the optimally transformed feature pairs in the agriculture data set (16 voivodeships and seven regions) in the system of the first two multiple kernel canonical variables $(U^{(1)},U^{(2)})$. The optimal transformations were determined by multiple kernel CCA for repeated measures data, as described in Section 2.
@@ 4.1.2. Multivariate functional data
We used a Fourier basis with nine components ($B_{ij}=9,\ i\in\{1,\ldots,L\},\ j\in\{1,\ldots,p_{i}\}$) to express the agriculture data as functional data. Using the transformed data, multiple functional canonical correlation analysis was performed (Table 1). The multivariate functional data for the 16 voivodeships in the system of the first two multiple functional canonical variables $(U^{(1)},U^{(2)})$ are shown in Figure 5, which shows that in the system of multiple canonical variables, the voivodeships are grouped into compact clusters belonging to specific macroregions. Additionally, looking at canonical correlations (Table 1), we see that higher values were obtained for the multivariate functional approach. Thus, for these data, multiple functional canonical variables have stronger discrimination ability than the multiple kernel canonical variables.
Figure 5. Scatterplot for the optimally transformed feature pairs in the agriculture data set (16 voivodeships and seven regions) in the system of the first two multiple functional canonical variables $(U^{(1)},U^{(2)})$. The optimal transformations were determined by multiple functional CCA, as described in Section 3.
@@ 12. Innovation 5
Figure 7. Scatterplot for the optimally transformed feature pairs in the GCI data set (115 countries and five groups) in the system of the first two multiple functional canonical variables $(U^{(1)},U^{(2)})$ (with 95% confidence normal ellipses). The optimal transformations were determined by multiple functional CCA, as described in Section 3.
@@ 5. Concluding remarks
The proposed operator-based framework for multiple CCA not only unifies existing approaches but also offers a principled way to handle dependence and non-compactness in high-dimensional or functional data. Future work may focus on extending these ideas toward regularized and sparse formulations, which could further improve stability and interpretability in complex empirical settings.